Managers use projected financial statements in four principal ways. (1) They use the projected statements to assess whether the firm's anticipated performance is in line with its own internal targets and with investors' expectations. (2) They use them to estimate the impact of proposed operating changes. (3) They use them to anticipate the firm's future financing needs and to arrange necessary financing. (4) Finally, they use them to estimate free cash flows, which determine the firm's overall value. Managers forecast free cash flows under different operating plans, forecast their capital requirements, and then choose the plan that maximizes shareholder value. Security analysts make the same types of projections as managers, and influence investors, who determine the future of firms' managers.

Increasing sales require additional assets, these assets must be financed, and it may or may not be possible to obtain all the funds needed for the firm's business plan. A key element in the financial forecasting process is to determine the (select one -internal or external) 1 of Item 1 financing requirements through the AFN equation. Additional funds needed are the amount of (select one -internal or external)2 of Item 1 capital (interest-bearing debt and preferred and common stock) that will be necessary to acquire the required assets. The AFN equation approximates the funds needed assuming that ratios (-Select one remain constant, fluctuate widely, only decline) of Item 1. The AFN equation is written as follows:

AFN = (A0*/S0)ΔS - (L0*/S0)ΔS - MS1(1 - Payout)

The AFN equation shows the relationship of -Select-internalexternalCorrect 4 of Item 1 funds needed by a firm to its projected increase in assets, the spontaneous increase in liabilities, and the increase in retained earnings. Rapidly growing companies require larger increases in assets; other things held constant, so -Select-expensesalesstockCorrect 5 of Item 1 growth is an important factor to the firm's AFN. The -Select-assets turnoverprice/earningscapital intensityCorrect 6 of Item 1 ratio is the ratio of assets required per dollar of sales. Companies with higher assets-to-sales ratios require more assets for a given increase in sales, hence a -Select-greatersmallerCorrect 7 of Item 1 need for external financing. -Select-Externally financedSpontaneously generatedBank obtainedCorrect 8 of Item 1 funds arise out of normal business operations from its suppliers, employees, and the government that reduce the firm's need for external financing. The -Select-lowerhigherCorrect 9 of Item 1 the profit margin, the larger the net income available to support increases in assets, hence the -Select-smallergreaterCorrect 10 of Item 1 the need for external financing. The -Select-payoutretentionquickCorrect 11 of Item 1 ratio is the proportion of net income that is reinvested in the firm, and it is calculated as 1 minus the -Select-P/E ratiodividend payout ratioquick ratioCorrect 12 of Item 1. The higher the -Select-payoutretentionquickCorrect 13 of Item 1 ratio, the lower the firm's AFN. The -Select-long-termsustainableexpectedGordonCorrect 14 of Item 1 growth rate is the maximum achievable growth rate without the firm having to raise external funds. In other words, it is the growth rate at which the firm's AFN equals zero.

Quantitative Problem 1: Beasley Industries' sales are expected to increase from $5 million in 2019 to $6 million in 2020, or by 20%. Its assets totaled $3 million at the end of 2019. Beasley is at full capacity, so its assets must grow in proportion to projected sales. At the end of 2019, current liabilities are $710,000, consisting of $170,000 of accounts payable, $400,000 of notes payable, and $140,000 of accrued liabilities. Its profit margin is forecasted to be 4%, and its dividend payout ratio is 50%. Using the AFN equation, forecast the additional funds Beasley will need for the coming year. Do not round intermediate calculations. Round your answer to the nearest dollar.
$  

The AFN equation assumes that ratios remain constant. However, firms are not always operating at full capacity so adjustments need to be made to the existing asset forecast. Excess capacity adjustments are changes made to the existing asset forecast because the firm is not operating at full capacity. For example, a firm may not be at full capacity with respect to its fixed assets. First, the firm's management must find out the firm's full capacity sales as follows:

Next, management would calculate the firm's target fixed assets ratio as follows:

Finally, management would use the target fixed assets ratio with the projected sales to calculate the firm's required level of fixed assets as follows:

Required level of fixed assets = (Target fixed assets/Sales) × Projected sales

Quantitative Problem 2: Mitchell Manufacturing Company has $1,700,000,000 in sales and $380,000,000 in fixed assets. Currently, the company's fixed assets are operating at 80% of capacity.

  1. What level of sales could Mitchell have obtained if it had been operating at full capacity? Do not round intermediate calculations. Round your answer to the nearest dollar.
    $  
  2. What is Mitchell's Target fixed assets/Sales ratio? Do not round intermediate calculations. Round your answer to two decimal places.
    %
  3. If Mitchell's sales increase by 40%, how large of an increase in fixed assets will the company need to meet its Target fixed assets/Sales ratio? Do not round intermediate calculations. Round your answer to the nearest dollar.
    $  
Answer #1

Quantitative Problem 1:

Solution :

Answer - 452,500$

Caluclations

AFN = (A0*/S0)ΔS - (L0*/S0)ΔS - MS1(1 - Payout)

A - Assets tied directly to sales

L- Sponteneous Liabilities that are affected by sales

S0 = The previous year sales

C = Total Projected sales for the next year

ΔS = Change of sales between S0 and S1

Ms1 = Net projected income

Retention Ratio = 1- payout

= 1-50%= 0.5

Assets = 3,000,000

Liabilities * 710,000

Expected projected sales (6,000,000 or 5,000,000+ (20%*5,000,000)

(25%*3,000,000)-(710000*25%)- 6,000,000*4%*(1-0.5)=

=750000-177500-120000

AFN =452,500$

Quantitative Problem 2:

Solution

a) Level of sales Mitchell could have obtained if it had been operating at full capacity = Current Sales/Current Operating Capacity

=  $1,700,000,000/80%

=2,125,000,000 $

b)Target fixed assets/Sales ratio

= $380,000,000/2,125,000,000

= 0.178 or 18% Rounding off to two decimal places

c) Target fixed assets/Sales ratio If Mitchell's sales increase by 40%

Projected sales = 1,700,000,000 +1,700,000,000*40%

=1,700,000,000+680,000,000=2,380,000,000

Required level of fixed assets = (Target fixed assets/Sales) × Projected sales

= 0.178*2,380,000,000

= 423,640,000

IF Mitchell's sales increase by 40%, increase in fixed assets will the company need to meet its Target fixed assets/Sales ratio = Required level of fixed assets - Current level of Fixed Assets

= 4,23,640,000-380000000

=43,640,000$